What is the slant height of a square pyramid that has a surface area of 525 square meters and a side length of 15 meters?
Answer by lwsshak3(11628)

What is the slant height of a square pyramid that has a surface area of 525 square meters and a side length of 15 meters?
...
A pyramid has 4 identical surfaces plus a square base
Area of base=15*15=225
Total area of surface=525-225=300
For one surface, area=300/4=75
area=1/2*base*height=75
1/2*15*h=75
7.5h=75
h=75/7.5=10 m